The ratio obtained from the coefficients in a balanced chemical equation is called the mole ratio. Use the balanced chemical reaction shown in model 2 to answer the following questions. 1. What is the mole ratio for the reaction in model 2 and explain why this is called a mole ratio. 2. If 0.052 mol N2 are reacted, how many mol NH3 are formed? Using dimensional analysis, show how you calculated your answer. 3. If 0.052 mol N2 are reacted, how many mol H2 are used? 4. If 0.863 mol NH3 are produced, how many mol N2 must have reacted? How many mol H2 must have reacted? 5 5. What mass of nitrogen is needed to produce 200.0 g of ammonia, NH3? Use the
Answer #1

1) The mole ratio is defined as the ratio between the number of moles of any two compounds involved in the chemical reaction

2) The balanced reaction is

N2 + 3H2 -------- 2NH3

one mole of N2 will yield 2 moles of NH3

Moles of NH3 formed = 2 * 0.052 = 0.104 moles of NH3

3) One mole of N2 reacts with 3 moles of H2

Moles of H2 reacted = 3 * 0.052 = 0.156 moles of H2

4) Number of moles of N2 reacted = 1/2 * number of moles of NH3 = 1/2 * 0.863 = 0.4315 moles

Number of moles of H2 reacted = 3/2 * number of moles of NH3 = 3/2 * 0.863 = 1.2945 moles

5) Moles of ammonia = Mass/Molar mass = 200/17 moles

Moles of N2 needed = 1/2 * Moles of ammonia = 100/17 moles

Mass of N2 needed = number of moles * molar mass = 100/17 * 28 = 164.7 grams

Note - Post any doubts/queries in comments section.